# vCache: Verified Semantic Prompt Caching

### v3

@@ 5 Evaluation
For our experiments, we use two popular embedding models – GteLargeENv1-5 (Zhang et al., 2024) and E5-large-v2 (Wang et al., 2022), and two LLM models for generating responses – Llama-3-8B-Instruct (Dubey et al., 2024) and GPT-4o-mini (gpt, 2024), representing both high-quality proprietary models and efficient open alternatives. We use the HNSW vector database (Malkov and Yashunin, 2018) with cosine similarity, a standard metric for comparing vector embeddings in semantic caching systems (Bang, 2023; Li et al., 2024; Dasgupta et al., 2025). All experiments are conducted on a machine running Ubuntu 24.04.2 LTS, equipped with an Intel Xeon Platinum 8570 CPU (for running cache) and an NVIDIA B200 GPU with 192 GB of memory (for running LLMs and embedding models).
@@
  • GPTCache (zilliztech, ): SOTA cache using a static threshold for all embeddings (Parameters: threshold).
  • GPTCache + Fine-tuned embedding (Zhu et al., 2024): Changing the embedding model in GPTCache. Embedding models are fine-tuned on data and method provided by (Zhu et al., 2024) for better separation of correct and incorrect matches (Parameters: threshold).
  • vCache: This is our proposed method (Parameters: error-rate guarantee $\delta$).
  • vCache + Fine-tuned embedding (Zhu et al., 2024): Same as vCache, but uses a fine-tuned embedding model (Parameters: error-rate guarantee $\delta$).
@@
To the best of our knowledge, no open-source benchmark currently exists for evaluating semantic caches. We introduce and open-source three diverse benchmarks designed to reflect common caching scenarios. Appendix D provides the complete dataset and benchmark cards.
@@
  • SemCacheClassification: A benchmark of 45,000 prompts derived from three classification datasets (Saurabh Shahane, 2023; Talmor et al., 2018; Ni et al., 2019), spanning 28 distinct labels. It simulates environments where LLMs are embedded in SQL queries to perform classification tasks (Liu et al., 2024b).
  • SemCacheSearchQueries: A random subset of 150,000 real-world queries from the ORCAS dataset (Craswell et al., 2020), consisting of short, search-style questions. It reflects scenarios where LLMs are used in web browsers to answer recurring user queries (Wang et al., 2024; Liu et al., 2024a), and caching might reduce redundant inference.
@@
We measure the following metrics. (1) Error Rate: (Lower is better) defined as the $FP/n$. where $n$ is the total number of prompts and FP is the number of false positives. (2) Cache hit rate: (Higher is better) defined as $(TP+FP)/n$ where TP and FP are true positives and false positives, respectively. Together, TP and FP measure the total number of cache hits. We also show ROC curves Hoo et al. (2017). For the non-deterministic evaluation of vCache (Algorithm 2, Line 5), we compute 95% confidence intervals using Wallis binomial confidence bounds and contingency tests Wallis (2013).
@@
We evaluate whether vCache satisfies the user-defined error rate $\delta$ while maintaining competitive performance. As shown in Figure 4 (left), vCache consistently meets the maximum error rate across $\delta$ values, with actual error remaining below the specified bound. The small gap between maximum error rate and observed error stems from the conservative $t^{\prime}$ estimation, which ensures robustness and can be further refined. Notably, as the error rate stabilizes, vCache continues to increase its cache hit rate (Figure 4, right), demonstrating effective learning over time. In contrast, GPTCache baselines show increasing error rates with sample size, reflecting the inherent limitations of fixed thresholds even as hit rates improve.
@@
We investigate whether semantic caches benefit from using dynamic, embedding-specific thresholds rather than a single static threshold. To evaluate this, we compare vCache against static-threshold baselines by systematically varying the threshold values for GPTCache-based baselines and the maximum error rate ($\delta$) for vCache-based caches across a feasible range. Each point in Figure 5 represents a complete evaluation over the full benchmark dataset, enabling a direct Pareto comparison between vCache and all static-threshold configurations. We find that vCache achieves better ROC curves, higher cache-hit rates at a given error rate, and consequently, significantly lower average latency. Specifically, in the SemCacheLMArena benchmark (Figure 5 top), vCache achieves up to 6× lower error, 2× higher cache hit rates, and reduced average latency. In SemCacheClassification (Figure 5 bottom), vCache outperforms all baselines for maximum error rates above approximately 1.5%, where static methods either violate the error constraint or underutilize the cache. For error rate bounds below 1.5%, vCache adopts a more conservative policy. This is a result of vCache’s modeling strategy, which prioritizes correctness by increasing exploration in regions of uncertainty.
Note that for a given threshold, GPTCache error rates have not yet converged and exhibit an increasing trend. Thus, the data presented here shows GPTCache in a more favorable light than it would appear had the error rates fully converged.
SemCacheSearchQueries Benchmark
The results for this benchmark are presented in Appendix C.1. We observe similar trends to those shown in Figure 4. vCache consistently meets the error guarantees set by the user, while the error rates of GPTCache continue to increase with more samples. In contrast to the previous two datasets, GPTCache’s error rates appear far from converging for SemCacheSearchQueries, potentially due to the high diversity of the dataset.
@@ 6 Limitations
There are two main limitations of vCache. First, for responses longer than a few words, vCache uses the LLM-as-a-judge approach (Zheng et al., 2023) to assess response equivalence (Algorithm 1, L8), which requires an additional inference call. However, this can be executed asynchronously outside the critical path, so it does not impact latency. Moreover, since the output is a single token (e.g., "yes" or "no"), the computational overhead is minimal (Leviathan et al., 2023). Second, vCache relies on the assumptions of i.i.d. data and a sigmoid function family to represent the probability of correctness. If these assumptions are violated, the analysis may not hold. Nonetheless, both assumptions are natural and appear to model most use cases well, as supported by our experimental results. Finally, while semantic caches introduce embedding overhead compared to LLM inference, this is generally offset by latency gains from cache hits, making the trade-off worthwhile on average (see Appendix C.2).
@@ 7 Conclusion
We introduced vCache, a semantic cache that provides correctness guarantees by learning an optimal similarity threshold for each cached embedding online. This approach addresses the limitations of static thresholds and embedding fine-tuning, and ensures that the error rate remains below a user-specified bound. Our experiments demonstrate that vCache consistently satisfies this guarantee while outperforming existing methods. These results suggest that reliable, interpretable caching for LLMs is both practical and deployable.
@@ Acknowledgement
The authors thank Sebastien Gautier-Jean-Jacques Beurnier and Siavash Ameli for their invaluable contributions. They also thank the Sky Lab at UC Berkeley for their support.
Appendix A Broader Impacts
This work contributes to the advancement of machine learning by improving the efficiency of large language model (LLM) inference. By reducing both computational cost and latency, our approach makes LLM-based systems more accessible to organizations and individuals with limited computational resources, thereby lowering the barrier to adoption. In addition, by decreasing the frequency of full LLM invocations, our method reduces overall compute demand and, consequently, the environmental impact associated with training and operating large-scale AI infrastructure.
Appendix B vCache Modeling Proof
@@ Appendix A Broader Impacts
Lemma B.1 (Upper-Bounding $\tau$).
@@
[…]
Lemma B.2 (Lower-Bounding Cache-Correctness Probability).
@@
$\begin{split}&\mathbf{Pr}(c(x)=1|\mathcal{D},x)\\ &=\mathbf{Pr}(t^{*}>t^{\prime}||\gamma^{*}<\gamma^{\prime})\cdot\mathbf{Pr}(c(x)=1|\mathcal{D},x,(t^{*}>t^{\prime}||\gamma^{*}<\gamma^{\prime}))\\ &\qquad+\mathbf{Pr}(\neg(t^{*}>t^{\prime}||\gamma^{*}<\gamma^{\prime}))\cdot\mathbf{Pr}(c(x)=1|\mathcal{D},x,t^{*}\leq t^{\prime},\gamma^{*}\geq\gamma^{\prime})\\ &\geq\mathbf{Pr}(\neg(t^{*}>t^{\prime}||\gamma^{*}<\gamma^{\prime}))\cdot\mathbf{Pr}(c(x)=1|\mathcal{D},x,t^{*}\leq t^{\prime},\gamma^{*}\geq\gamma^{\prime})\\ &\geq\;(1-\epsilon)\,\cdot\mathbf{Pr}\bigl{(}c(x)=1\mid\mathcal{D},x,t^{*}\leq t^{\prime},\gamma^{*}\geq\gamma^{\prime}\bigr{)}.\\ \end{split}$ (19)
@@
To find $t^{\prime},\gamma^{\prime}$ from the estimated $\hat{t},\hat{\gamma}$, we can use the confidence intervals by assuming a uniform prior on $t^{*}$ and $\gamma^{*}$. Since, under uniform prior the distributions of $\mathbf{Pr}(t^{*}|\hat{t})$ and $\mathbf{Pr}(\hat{t},t^{*})$ are the same,
@@
Appendix C Evaluation Results
C.1 SemCacheSearchQueries Benchmark
@@
Figure 7: Pareto comparison across a range of thresholds and $\delta$ values on 150k samples from the SemCacheSearchQueries benchmark. While vCache slightly underperforms static baselines at this scale, its error rate remains bounded. In contrast, static-threshold methods like GPTCache require increasing thresholds to maintain low error, shifting the Pareto curve upward over time. This highlights the limitations of static thresholds in sustaining error-rate guarantees as sample size grows.
@@
We evaluate whether dynamic, embedding-specific thresholds yield better long-term performance than static thresholding. To this end, we compare vCache against GPTCache by varying static similarity thresholds and vCache’s maximum error rate bound $\delta$. Each point in Figure 7 reflects a full evaluation over the SemCacheSearchQueries benchmark, enabling a direct Pareto comparison. Static-threshold configurations achieve higher cache hit rates and lower latencies than vCache under equivalent evaluation settings. However, this performance comes at the cost of increasing error as the sample size grows, as shown in Figure 6. Because the threshold remains fixed, both hit rate and error rate rise together. As a result, the GPTCache curve in the middle plot (cache hit vs. error rate) is expected to shift up and to the right, while the curve in the right plot (error rate vs. latency) shifts up and to the left. This trend implies that no static threshold below 1.0 can maintain a bounded error rate as prompt diversity increases. In contrast, vCache learns its threshold online and per embedding, allowing it to enforce the error constraint while gradually improving cache hit rate. While this results in a more conservative reuse policy initially, vCache offers greater reliability under error-sensitive deployment conditions.
C.2 Embedding Generation Overhead
@@
Semantic caches incur additional latency due to embedding computation, which must be evaluated in relation to the cost of LLM inference. To quantify this overhead, we compare the embedding latencies of four models—including fine-tuned variants—to the inference latency of Llama3.1-8B.
@@
Figure 8: Embedding computation latency distributions across models, shown as 95th percentile whisker plots. GTE_FT = GteLargeENv1-5 [Zhang et al., 2024] fine-tuned [Zhu et al., 2024]. E5 Large v2 FT = E5-large-v2 [Wang et al., 2022] fine-tuned [Zhu et al., 2024]
@@
All experiments are conducted on a system with an Intel Xeon Platinum 8570 CPU and an NVIDIA B200 GPU with 192 GB of memory. The results show that embedding computation is significantly faster than model inference, confirming its applicability in caching pipelines.
C.3 Threshold Dilemma

### v4

@@ 5 Evaluation
For our experiments, we use two popular embedding models (GteLargeENv1-5 (Zhang et al., 2024) and E5-large-v2 (Wang et al., 2022)), and two LLMs (Llama-3-8B-Instruct (Dubey et al., 2024) and GPT-4o-mini (gpt, 2024)), representing both high-quality proprietary models and efficient open alternatives. We use the HNSW vector database (Malkov & Yashunin, 2018) with cosine similarity, a standard metric for comparing vector embeddings in semantic caching systems (Bang, 2023; Li et al., 2024; Dasgupta et al., 2025). All experiments are conducted on a machine running Ubuntu 24.04.2 LTS, with an Intel Xeon Platinum 8570 CPU and an NVIDIA Blackwell with 192 GB of memory.
@@
  • GPTCache (zilliztech, ): SOTA using a static threshold for all embeddings (Parameter: threshold).
  • GPTCache + Fine-tuned embedding: Changing the embedding model in GPTCache. Embedding models are fine-tuned on data and method provided by Zhu et al. (2024) (Parameter: threshold).
  • vCache: This is our proposed method (Parameter: error-rate bound $\delta$).
  • vCache + Fine-tuned embedding (Zhu et al., 2024): Same as vCache, but uses a fine-tuned embedding model (Parameter: error-rate bound $\delta$).
@@
To the best of our knowledge, no realistic open-source benchmark currently exists for evaluating semantic caches. We introduce and open-source three diverse benchmarks designed to reflect common caching scenarios. Appendix E provides the complete dataset and benchmark cards.
@@
  • SemCacheClassification: A benchmark of 45,000 prompts derived from three classification datasets (Saurabh Shahane, 2023; Talmor et al., 2018; Ni et al., 2019). Simulating environments where LLMs are embedded in SQL queries to perform classification tasks (Liu et al., 2024b).
  • SemCacheSearchQueries: A random subset of 150,000 real-world queries from the ORCAS dataset (Craswell et al., 2020), consisting of short, search-style questions. It reflects scenarios where LLMs are used in web browsers to answer recurring user queries (Wang et al., 2024; Liu et al., 2024a), and caching might reduce redundant inference. The results are presented in Appendix D.1.
@@
We measure the following metrics. (1) Error Rate: (Lower is better) defined as $FP/n$, where FP is the number of false positives and $n$ is the total number of prompts. (2) Cache hit rate: (Higher is better) defined as $(TP+FP)/n$ where TP and FP are true positives and false positives, respectively. Together, TP and FP measure the total cache hits. We also show ROC curves Hoo et al. (2017). For the non-deterministic evaluation of vCache (Algorithm 2, Line 5), we compute 95% confidence intervals using Wallis binomial confidence bounds and contingency tests Wallis (2013).
@@
We evaluate whether vCache satisfies the user-defined error rate bound $\delta$ while maintaining competitive performance. As shown in Figure 4 (left), vCache consistently meets the maximum error rate across $\delta$ values, with actual error remaining below the specified bound. The small gap between maximum error rate and observed error stems from the conservative $t^{\prime}$ estimation, which ensures robustness and can be further refined. Notably, as the error rate stabilizes, vCache continues to increase its cache hit rate (Figure 4, right), demonstrating effective learning over time. In contrast, GPTCache baselines exhibit increasing error rates with sample size, reflecting the inherent limitations of fixed thresholds despite improving hit rates.
@@
We evaluate whether semantic caches benefit from dynamic, embedding-specific thresholds over a single static threshold. To this end, we compare vCache with static-threshold baselines by systematically varying threshold values for GPTCache and the maximum error rate ($\delta$) for vCache across a feasible range. Each point in Figure 5 represents a complete evaluation over the benchmark dataset, enabling direct Pareto comparison between vCache and static-threshold configurations. vCache achieves better ROC curves, higher cache-hit rates at a given error rate, and lower average latency. On SemCacheLMArena (Figure 5, top), it achieves up to 6× lower error, 2× higher hit rates, and reduced latency. On SemCacheClassification (Figure 5, bottom), vCache outperforms all baselines for error bounds above  1.5%, where static methods either violate constraints or underutilize the cache. For bounds below 1.5%, vCache is more conservative, reflecting its strategy of prioritizing correctness by increasing exploration under uncertainty. We note that GPTCache error rates have not fully converged and exhibit an upward trend, suggesting the reported results likely overstate GPTCache performance.
Latency, vLLM, and Additional Baselines
We report additional experiments on embedding generation overhead (Appendix D.2), logistic regression latency overhead (Appendix D.3), vCache performance combined with vLLM (Appendix D.4), and extended baseline evaluations (Appendix D.5).
@@ 6 Limitations
There are two main limitations of vCache. First, for responses longer than a few words, string matching is insufficient and vCache uses a standard LLM-as-a-judge (Zheng et al., 2023a) approach to assess response equivalence (Algorithm 1, L8), requiring an additional LLM inference. However, this can be executed asynchronously outside the critical path, so it does not impact latency (see vCache implementation). Moreover, since the output is a single token (e.g., ”yes” or ”no”), the computational overhead is minimal (Leviathan et al., 2023). In SemCacheLMArena and SemCacheSearchQueries, response equivalence is assessed with LLM-as-a-judge, whereas in SemCacheClassification it is determined by string matching. vCache performs reliably under both evaluation regimes. Second, vCache relies on the assumptions of independent and identically distributed (i.i.d.) data and a sigmoid function family to represent the probability of correctness. If these assumptions are violated, the analysis may not hold. Nonetheless, both assumptions are natural and appear to model most use cases well, as supported by our experimental results.
@@ 7 Conclusion
We introduced vCache, a semantic cache that provides correctness guarantees by learning an optimal similarity threshold for each cached embedding online. This approach addresses the limitations of static thresholds and embedding fine-tuning in the semantic caching domain, ensuring that the error rate remains below a user-specified bound. Our experiments demonstrate that vCache consistently satisfies this guarantee while outperforming existing methods. These results suggest that reliable, interpretable caching for LLMs is both practical and deployable.
@@ Acknowledgement
The authors thank Sebastien Gautier-Jean-Jacques Beurnier and Siavash Ameli for their invaluable contributions. They also thank the Sky Computing Lab at UC Berkeley for their support.
Appendix A Notation Glossary
For clarity, we summarize the key symbols and functions used throughout the paper in Table 1. The glossary covers cache contents, decision policies, probability functions, and modeling parameters. Each entry is accompanied by a short explanation and, where applicable, references to the defining equations, algorithms, or sections.
Table 1: A glossary of notations used in the paper and their explanations
Notation Explanation
$r(x)$ Response produced by the LLM for prompt $x$.
$\mathcal{E}(x)$ Vector embedding of prompt $x$.
$\mathrm{nn}(x)$ Current nearest neighbor of $x$ in the cache; may change as the cache grows (see Equation 2).
$s(x)$ Similarity between prompt $x$ and $\mathrm{nn}(x)$ (see Equation 4).
$c(x)$ Correctness: $1$ if $r(\mathrm{nn}(x))=r(x)$, else 0 (see Equation 4).
$\mathcal{D}$ Cache contents: $(\mathcal{E}(x_{i}),r(x_{i}),\mathcal{O}(x_{i}))_{i=1}^{n}$ (see Equation 3).
$\mathcal{O}(x_{i})$ Metadata for cached prompt $x_{i}$: all observed $(s(x_{j}),c(x_{j}))$ with $\mathrm{nn}(x_{j})=x_{i}$ (see Equation 3).
$\mathcal{P}_{\text{gptCache}}(s(x))$ Static-threshold policy used by existing systems.
$\mathcal{P}_{\text{vCache}}(s(x),\mathcal{O}(\mathrm{nn}(x)),\delta)$ vCache policy using embedding-specific modeling under a user-defined error bound $\delta$ (Section 4).
$\delta$ User-defined maximum error tolerance (see Section 4.1).
$\mathrm{Pr}(\mathrm{vCache}(x)=r(x))$ Probability that vCache returns the correct response (see Equation 8).
$\tau_{\mathrm{nn}(x)}(s(x))$ Required exploration probability to satisfy the guarantee (see Equation 8).
$\hat{\tau}$ Upper bound used for exploration probability (see Equation 11).
$u\sim\mathrm{Uniform}(0,1)$ Random draw used to realize the exploration decision (see Algorithm 2).
$\mathcal{L}(s(x),t,\gamma)$ Sigmoid likelihood modeling $\mathbf{Pr}(c(x)=1\mid x,\mathcal{D})$ (see Equation 9).
$t,\gamma$ True threshold ($t$) and slope ($\gamma$) parameters of $\mathcal{L}$ (see Equation 9).
$\hat{t},\hat{\gamma}$ MLE estimates of $t$ and $\gamma$ based on $\mathcal{O}(\mathrm{nn}(x))$ (see Equation 10).
$t^{\prime},\gamma^{\prime}$ Conservative estimates of the logistic parameters $t,\gamma$, selected from the confidence band of the MLE estimates. These values are used to compute $\hat{\tau}$ and ensure that vCache respects the user-defined error bound $\delta$ (see Equation 11).
Appendix B vCache Semantic Cache Architecture
Figure 6 illustrates the vCache architecture. When a new prompt gets processed, it is first embedded into a vector representation and queried against the vector database to retrieve the nearest cached prompt. The similarity score and metadata of the retrieved embedding are passed to the similarity evaluator, which compares the correctness estimate against the user-defined error bound $\delta$ (see vCache decision policy in Section 4.2). If the policy returns $\mathrm{exploit}$, the cached response is retrieved from the response store and served immediately. If the policy returns $\mathrm{explore}$, the system performs an LLM inference to generate the true response, determines the correctness of the cached response with respect to the newly generated one, updates the metadata of the nearest neighbor, adds the new embedding to the vector database, adds the generated response to the response store, and returns it to the user (see Algorithm 1). In the vector database, the green balls represent the confidence bounds specific to the currently processed prompt. Larger balls indicate lower thresholds (more conservative exploitation). In comparison, smaller balls correspond to higher thresholds (more conservative exploration).
Figure 6: Workflow of the vCache architecture. Prompts are embedded, nearest neighbors retrieved, and the decision policy selects between exploiting a cached response or exploring via an LLM inference while ensuring the user-defined error bound $\delta$.
Appendix C vCache Modeling Proof
@@ Appendix A Notation Glossary
Lemma C.1 (Upper-Bounding $\tau$).
@@
$\mathbf{Pr_{explore}}(x,\mathcal{D}_{n})\;\geq\;\tau^{\prime}\;=\;1\;-\;\frac{\delta}{1-\alpha}\geq\;\tau_{\mathrm{nn}(x)}(s(x)).$ (15)
Lemma C.2 (Lower‐Bounding Cache‐Correctness Probability).
@@
[…]
@@
[…]
@@
Appendix D Evaluation Results
D.1 SemCacheSearchQueries Benchmark
@@
Figure 8: Pareto comparison across a range of thresholds and $\delta$ values on 150k samples from the SemCacheSearchQueries benchmark. While vCache outperforms static baselines while its error rate remains bounded. In contrast, static-threshold methods like GPTCache require increasing thresholds to maintain low error (i.e., a threshold of 0.99 for a 1.7% error rate after 150k samples), shifting the Pareto curve upward over time. This highlights the limitations of static thresholds in sustaining error-rate guarantees as sample size grows.
@@
We evaluate whether dynamic, embedding-specific thresholds yield better long-term performance than static thresholding. To this end, we compare vCache against GPTCache by varying static similarity thresholds and vCache’s maximum error rate bound $\delta$. Each point in Figure 8 reflects a complete evaluation over the SemCacheSearchQueries benchmark, enabling a direct Pareto comparison. Static-threshold configurations achieve lower cache hit rates and higher error rates than vCache under equivalent evaluation settings. As shown in Figure 7, GPTCache has an increasing error rate as the sample size grows because the threshold remains fixed (both cache hit rate and error rate rise together). As a result, the GPTCache curve in the middle plot (cache hit vs. error rate) is expected to shift up and to the right, while the curve in the right plot (error rate vs. latency) shifts up and to the left. This trend suggests that no static threshold below 1.0 can maintain a bounded error rate as prompt diversity increases (see the threshold of 0.99, which yields an error rate of 1.7% after 150k samples). In contrast, vCache learns its threshold online and per embedding, allowing it to enforce the error constraint while gradually improving cache hit rate.
D.2 Embedding Generation Overhead
@@
Semantic caches incur additional latency due to embedding computation, which must be evaluated in relation to the cost of LLM inference. To quantify this overhead, we compare the embedding latencies of four models to the inference latency of Llama3.1-8B.
@@
Figure 9: Embedding computation latency distributions across models, shown as 95th percentile whisker plots. GTE_FT = GteLargeENv1-5 (Zhang et al., 2024) fine-tuned (Zhu et al., 2024). E5 Large v2 FT = E5-large-v2 (Wang et al., 2022) fine-tuned (Zhu et al., 2024)
@@
All experiments are conducted on a system with an Intel Xeon Platinum 8570 CPU and an NVIDIA Blackwell GPU with 192 GB of memory. The results show that embedding computation is significantly faster than model inference, confirming its applicability in caching pipelines.
D.3 Logistic Regression Latency Overhead
Since vCache performs threshold estimation online for every request, the efficiency of logistic regression directly impacts scalability. In our experiments, we use sklearn.linear_model on CPU, yielding an average latency of 0.0017 seconds on the SemBenchmarkArena. Its negligible latency ensures that online modeling does not introduce noticeable overhead and can scale to large caches.
D.4 vCache in combination with vLLM
vCache is orthogonal to inference optimization systems, as semantic prompt caching reuses responses for semantically similar prompts rather than accelerating inference itself. When a cache miss occurs, vCache directly benefits from systems such as vLLM (Kwon et al., 2023) or SGLang (Zheng et al., 2023b), which reduce model latency. To validate this, we hosted LLaMa 3.1 70B with vLLM on two NVIDIA Blackwell GPUs and compared inference latency with and without vCache. For vCache, we additionally ran the GteLargeENv1_5 embedding model on the same machine. Evaluation was conducted on 45k samples from the SemCacheClassification benchmark.
Table 2: A comparison of overall runtime, latency, cache hit rate, and error rate with and without vCache under different error tolerances
Baseline Config
Overall
Duration
Avg. LLM
Inference Latency
Avg. Emb.
Latency
Cache
Hit Rate
Error
Rate
vLLM 240 min 0.32 sec 0.0% 0.0%
vLLM
+ vCache
$\delta=0.01$ 214 min 0.32 sec 0.018 sec 18.1% 0.4%
$\delta=0.02$ 170 min 0.32 sec 0.018 sec 35.5% 1.2%
$\delta=0.03$ 160 min 0.32 sec 0.018 sec 40.2% 1.4%
Despite the additional embedding overhead, vCache substantially reduces end-to-end latency by avoiding repeated LLM inferences. This demonstrates that vCache is complementary to inference optimization systems such as vLLM.
D.5 Additional Baseline Evaluations
vCache is the first adaptive, probabilistic, and Bayesian method for semantic caching. The most competitive alternatives are static threshold methods, which we extend with several additional baselines for completeness. The landscape of approaches can be ordered from naive to advanced as follows:
GS:
Global and Static threshold (i.e., GPTCache).
GD:
Global and Dynamic threshold (vCache with global threshold).
LS:
Per-embedding (Local) and Static threshold.
LD:
Per-embedding (Local) and Dynamic threshold:
LD1:
Logistic regression to compute the threshold ($\hat{t}$).
LD2:
Logistic regression sigmoid fit to model correctness probability.
LD3:
Logistic regression sigmoid fit with confidence intervals and guarantees (vCache).
We implement all baselines and perform ablations on the E5-large-v2 embedding model, OpenAI GPT-4o mini LLM, and the SemCacheLmArena benchmark. For non-adaptive baselines (LD1 and LD2), we select the threshold or $\delta$ values that produced error rates closest to their observed performance. Results are shown in Table 3.
Table 3: A comparison of error rate, cache hit rate, and qualitative observations across baselines
Baseline
Threshold/
Delta
Error
Rate
Cache hit
Rate
Comments and Observations
GS 0.99 2.5% 37% No guarantee and worst trade-off.
0.98 4.1% 53%
0.97 5.2% 67%
GD 0.02 1.3% 14% Due to the large overlap between incorrect and correct samples at a given similarity (see 3), the optimal threshold converges to 1.0, yielding low cache hits.
0.03 2.5% 26%
0.05 4.3% 45%
LS Impossible to compute a threshold for every embedding a priori.
LD1 2.6% 70% No guarantees and no error-rate fine-tuning.
LD2 2.1% 68% No guarantees and no error-rate fine-tuning.
LD3 0.02 0.5% 41% Guarantees and beats both SOTA and GD baselines.
0.03 1.1% 46%
0.05 2.0% 54%
This ablation underlines that (1) semantic caches benefit from embedding-specific thresholds, and (2) probabilistic modeling is required to satisfy user-defined error bounds and ensure predictability. Notably, only vCache provides guarantees. The importance of guarantees cannot be overstated, as in practice, the absence of correctness guarantees has been a primary reason for the failure of semantic caching deployments in industry.
D.6 Threshold Dilemma